OpenStudy (anonymous):

Real quick, last one! Find the x-intercepts of the parabola with the vertex (-1,2) and y-intercept (0,-3)

myininaya (myininaya):

\[f(x)=a(x-h)^2+k\] where (h,k) is vertex but we vertex is (-1,2) => h=-1 and k=2 so we have \[f(x)=a(x-(-1))^2+2\] now we are given f(0)=-3 use this to find a

OpenStudy (anonymous):

???

OpenStudy (zarkon):

Solve \[-3=a(0+1)^2+2\]for a

OpenStudy (anonymous):

Ok hang on

OpenStudy (zarkon):

that will give you the quadratic... then you can find the roots.

OpenStudy (anonymous):

Sorry I am completely awful at looking at something and turning it into the problem I need

OpenStudy (anonymous):

a=-5

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Where do I plug that in?

OpenStudy (zarkon):

into \[f(x)=a(x+1)^2+2\]

OpenStudy (anonymous):

Ok hang on

OpenStudy (zarkon):

then solve \[f(x)=0\] for x

OpenStudy (anonymous):

What do I do to f(x)=-5(x+1)^2+2 find the vertex or what?

OpenStudy (zarkon):

you want the x-intercepts so solve f(x)=0 for x

OpenStudy (zarkon):

solve \[-5(x+1)^2+2=0\]

OpenStudy (anonymous):

0.37 and -1.63 is what I got for the x intercepts...

OpenStudy (anonymous):

is that correct?

OpenStudy (zarkon):

-.37

OpenStudy (zarkon):

\[-1\pm\frac{\sqrt{10}}{5}\]

OpenStudy (zarkon):

(-.367544467966,0) (-1.63245553203,0)

OpenStudy (anonymous):

Thank you so much for helping me

OpenStudy (zarkon):

np